OpenStudy (anonymous):

Find a particular solution to y"+25y=20sin(5t). I figured that Y = Asin(5t), so Y' = 5Acos(5t) and Y" = -25Asin(5t), but if you put Y and Y" in the equation, you get 0 = 20sin(5t)... so do you add a t to the initial Y?

OpenStudy (freckles):

hey

OpenStudy (freckles):

weren't you the one I showed that site to earlier?

OpenStudy (freckles):

if have the right side is sin( t) or cos(t ) or sin(t)+cos(t) then the particular solution guess should be y=Asin(t)+Bcos(t) so here you should have y=Asin(5t)+Bcos(5t)

OpenStudy (anonymous):

Yes, and I've checked it but it really only has a few examples, such as y" - 4y' -12y = sin(2t).

OpenStudy (anonymous):

Interesting... I wish my professor had gone more in depth with this type.

OpenStudy (freckles):

the table should say this though

OpenStudy (freckles):

and it does on row 3

OpenStudy (anonymous):

Ah here we go, my table says \[e^(\alpha*t) (Q_n *t sinBt + R_ncosBt)\]. I didn't see all of it I suppose. Thanks!

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx I was talking about row 3 here: g(t)=20sin(5t) so y_p=Acos(5t)+Bsin(5t)

ganeshie8 (ganeshie8):

I think, we can get the right hand side by differentiating either `sin` or `cosine` multiple times, so your guess must be a combination of both these functions

ganeshie8 (ganeshie8):

when the right hand side is sin or cosine or an exponential, we can express it as \(e^{ax}\) and then there is an easy direct formula for the particular solution : \[y_p = \dfrac{e^{ax}}{p(a)}\]

ganeshie8 (ganeshie8):

don't bother about it if your professor haven't covered them yet..

OpenStudy (anonymous):

It seems that it still comes out to 0. Y" = -25Asin(5t) -25Bcos(5t), Y = Asin(5t)+Bcos(5t). Multiply Y by 25, -25+25 = 0.

ganeshie8 (ganeshie8):

whats your homogeneous solution ?

OpenStudy (anonymous):

ganeshie8, what does p(a) signify then?

OpenStudy (freckles):

oh you are right

OpenStudy (freckles):

so we have to adjust the particular solution

OpenStudy (freckles):

I think (thinking)

OpenStudy (anonymous):

So then maybe Y = At^2 sin(5t) + Bt^2cos(5t)?

ganeshie8 (ganeshie8):

\(y'' + 25y\) can be expressed as \((D^2 + 25)y\) \(p(D) = D^2+25\) Hey don't bother about these, it is better you get these in class from your prof

ganeshie8 (ganeshie8):

lets focus on undetermined coefficients :)

OpenStudy (anonymous):

Our homework is due tonight and he assigned these problems.

ganeshie8 (ganeshie8):

Do you know why you're stuck ?

ganeshie8 (ganeshie8):

Find the homogeneous solution and you will find the reason

OpenStudy (anonymous):

Homogeneous as in the \[y_H\]?

OpenStudy (freckles):

oh... we need to multiply t to the particular guess we had right @ganeshie8 ?

OpenStudy (anonymous):

Ah, so our c1 and c2 are imaginary numbers: +/- 5i.

OpenStudy (anonymous):

So \[y_H = c_1e^(5t) + c_2e^(5t)?\]

ganeshie8 (ganeshie8):

when you have complex roots, \(r=\pm 5i\), the homogeneous solutoin can be written as : \[y_H = c_1\cos(5t) + c_2\sin(5t)\]

ganeshie8 (ganeshie8):

Notice that your previous guess for particular solution is actually the solution to the homogeneous equaiton. So it never gonna work.

ganeshie8 (ganeshie8):

just multiiply `t` to the previous guess as freckles is saying and see...

OpenStudy (anonymous):

I think I figured it out with multiplying by t, I just went back to the singular Y = Atsin(5t), and so my answer came out to A = 2tan(5t).

OpenStudy (anonymous):

Still not coming up with the right answer though, my Y would then be \[2t \tan(5t)*\sin(5t)\]

ganeshie8 (ganeshie8):

thats the solution generated by wolfram using undetermined coefficients you will be able to zoom in.. resolution is not an issue

OpenStudy (anonymous):

Wow, wolfram alpha really does its job. I'll have to go over that and figure it out through the steps. Thanks very much!

ganeshie8 (ganeshie8):

the first half has steps to find the homogeneous solution the second half is about the particular solution

ganeshie8 (ganeshie8):

the steps it has taken are exactly same as yours

OpenStudy (anonymous):

Oh man, I had the exact same thing when I did Y = Atsin(5t) +Btcos(t)... I just didn't simplify.

OpenStudy (anonymous):

Btcos(5t)*

ganeshie8 (ganeshie8):

I see.. happens haha!

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

It gives the exact reason for why we need to multiply our guess by "t" too... thats the reason i like solutions generated by wolfram but unfortunately it fails to generate solutions if the problems are a bit complicated..

OpenStudy (anonymous):

Ah, I understand now. Some of this stuff is just small mistakes, and the problems take a pretty long time I suppose. Thanks again!

ganeshie8 (ganeshie8):

np